# OminiControl: Minimal and Universal Control for Diffusion Transformer

### v2

@@ 3.4 Condition strength factor
$\text{bias}(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (7)
This design ensures that the strength factor $\gamma$ only affects the attention weights between noisy image tokens and condition image tokens, while maintaining the original attention patterns for text tokens and within-modality interactions.
3.5 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, vehicles, and animals, totaling over 200,000 images. This dataset, along with the generation pipeline, will be publicly released.
Training models for subject-consistent generation typically requires paired images that maintain identity consistency while exhibiting variations in pose, lighting, and other attributes. Previous methods like IP-Adapter [39] use identical images for conditioning and target pairs, which proves effective for their approaches. However, in our framework, this setup leads to overfitting, causing the model to generate outputs nearly identical to the inputs.
To overcome these limitations, we developed a dataset featuring images that preserve subject identity while incorporating natural variations. While existing datasets [31, 12, 15, 17] address similar needs, they often face constraints in either quality or scale. We therefore propose a novel synthesis pipeline leveraging FLUX’s inherent capability to generate pairs of visually related images from carefully crafted prompts.
Our pipeline utilizes ChatGPT-4o to generate over 20,000 diverse image descriptions, which guide FLUX in producing more than 200,000 images (Figure 6). The generated images undergo quality assessment using ChatGPT-4o’s vision capabilities, ensuring both consistency and diversity in the final dataset. The curated dataset and complete generation pipeline are publicly available [^3] More details are provided in the supplementary material. Dataset and code are available at https://github.com/Yuanshi9815/Subjects200K.
@@ 4.1 Setup
[…]
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.75
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.74
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.53
Ours 0.38 20.63 0.40 0.61 75.91 0.76
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.726
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.72
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.54
Ours 903 27.26 0.39 0.55 75.06 0.728
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.78
Ours 132 11.49 0.87 0.39 67.63 0.87
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.85
Ours 24 10.23 0.73 0.43 70.74 0.90
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.84
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.80
@@
Training. Our model is trained with a batch size of 1 and gradient accumulation over 8 steps (effective batch size of 8). We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01. The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each). For spatially aligned tasks, models are trained for 50,000 iterations, while subject-driven generation models are trained for 15,000 iterations.
Baselines. For spatially-aligned tasks, we compare our method with both the original ControlNet [41] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [14], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [39], evaluating its implementations FLUX.1 [37].
@@
Evaluation protocol. We conducted evaluations on two datasets. For spatially-aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
@@ 4.3 Empirical studies
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like depth-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.

### v4

@@ 3.3 Subjects200K datasets
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.304
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.305
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.192
Ours 0.38 20.63 0.40 0.61 75.91 0.308
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.308
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.309
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.212
Ours 903 27.26 0.39 0.55 75.06 0.308
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.293
Ours 132 11.49 0.87 0.39 67.63 0.300
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.278
Ours 24 10.23 0.73 0.43 70.74 0.301
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.300
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.306
To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [15] can generate identity-consistent image pairs when provided with appropriate prompts. Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^3] Detailed prompt design and synthesization pipeline are provided in the Appendix A.:
  • Prompt Generation. We use GPT-4o to generate over 30,000 diverse subject descriptions. Each description represents the same subject in multiple scenes. The goal is to highlight distinct objects or subjects while allowing natural variations.
  • Paired-image synthesis. We reorganized the collected descriptions into structured prompts. Each prompt describes the same subject in two different scenes. The template of such prompt is shown in Supplementary Material S3.These prompts are then fed into FLUX to generate image pairs. Each pair is designed to maintain subject consistency while varying in pose, lighting, and context.
  • Quality Assesment. Finally, we use GPT-4o to evaluate the generated pairs. Misaligned pairs are removed to ensure identity consistency and high image quality.
The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories including clothing, furniture, vehicles, and animals. Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control. To facilitate future research, we release both the dataset and our complete generation pipeline [^4] Dataset and code available at: https://github.com/Yuanshi9815/Subjects200K..
@@ 4.1 Setup
Training. Our model is trained with a batch size of 1 and gradient accumulation over 8 steps (effective batch size of 8). We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01. For spatially aligned tasks, we use text-to-image-2M[8] dataset with the last 300,000 images. For subject-driven generation, we utilize our proposed Subjects200K dataset. The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each). For spatially aligned tasks, models are trained for 50,000 iterations, while subject-driven generation models are trained for 15,000 iterations.
Baselines. For spatially aligned tasks, we compare our method with both the original ControlNet [42] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [15], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [41], evaluating its implementations FLUX.1 [39].
@@
Evaluation protocol. We conducted evaluations on two datasets. For spatially aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
@@ 4.3 Empirical studies
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and spatially non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.